# IDT: A Physically Grounded Transformer for Feed-Forward Multi-View Intrinsic Decomposition

### v1

@@ 4.5 Quantitative Results
Quantitative results on Hypersim are reported in Table 1. IDT consistently outperforms single-view intrinsic decomposition baselines in terms of intrinsic accuracy while achieving substantially better multi-view consistency. Compared to the feed-forward VGGT per-view baseline, IDT yields significantly lower cross-view inconsistency for both albedo and diffuse shading, while maintaining comparable or improved reconstruction quality.
These results demonstrate that joint multi-view inference with shared intrinsic representations is critical for enforcing physically meaningful consistency across views.
We further report explicit multi-view consistency metrics in Table 2, where IDT achieves the lowest inconsistency across all evaluated methods.
@@ 4.6 Ablation Studies
We conduct ablation studies on Hypersim to analyze the contribution of key components in IDT.
Effect of joint multi-view inference.
We compare IDT with a per-view variant that processes each view independently. Removing joint multi-view aggregation results in a clear degradation of cross-view consistency, confirming the importance of transformer-based joint reasoning.
Effect of intrinsic adapters.
We evaluate a variant of IDT without intrinsic adapters, where prediction heads directly consume aggregated transformer tokens. This variant exhibits reduced intrinsic decomposition quality and increased entanglement between material and illumination factors, highlighting the role of adapters in guiding factor-specific prediction.
Effect of illumination conditioning.
We further evaluate a variant of IDT without SGM-based illumination conditioning. Removing illumination conditioning leads to less stable shading estimates and increased leakage of illumination effects into albedo, demonstrating the importance of a shared illumination representation even without explicit physically based rendering.
@@ 4.7 Qualitative Results
Figure 4 presents qualitative comparisons on Hypersim and InteriorVerse. On Hypersim, IDT produces cleaner diffuse reflectance and more coherent diffuse shading compared to baseline methods, closely matching ground truth intrinsic factors. On InteriorVerse, IDT generalizes well to realistic indoor scenes, yielding view-consistent albedo and stable shading across viewpoints. Notably, specular highlights are better isolated into the specular shading component, reducing contamination of diffuse reflectance.
Overall, qualitative results further confirm the advantages of physically grounded, feed-forward multi-view intrinsic decomposition.
@@ 5 Conclusion
We presented Intrinsic Decomposition Transformer (IDT), a feed-forward framework for physically grounded multi-view intrinsic image decomposition. By jointly reasoning over multiple views with a transformer architecture, IDT infers view-consistent intrinsic factors in a single forward pass, avoiding iterative generative inference. Through an explicit image formation model that separates diffuse reflectance, diffuse shading, and specular effects, the proposed formulation improves the disentanglement of material and illumination across views. Experiments on both synthetic and real-world indoor datasets demonstrate that IDT yields cleaner intrinsic decompositions and significantly improved multi-view consistency compared to prior methods. We believe IDT provides a simple yet effective foundation for scalable multi-view intrinsic understanding and can serve as a building block for physically grounded reasoning in downstream vision tasks.

### v2

@@ 4.5 Quantitative Results
Single-view results.
Single-view quantitative comparisons are reported in Tables 2, 3, and 1. IDT consistently matches or outperforms state-of-the-art methods on depth, surface normals, and intrinsic decomposition, demonstrating that joint modeling does not compromise per-view accuracy.
@@ 4.6 Ablation Studies
We conduct ablation studies on Hypersim to analyze key components of IDT.
Joint multi-view inference.
Removing joint multi-view attention and processing each view independently leads to a significant degradation in cross-view consistency, while single-view accuracy remains comparable.
Intrinsic adapters.
Disabling intrinsic-specific adapters results in increased entanglement between material and illumination factors, leading to degraded albedo and shading quality.
Illumination conditioning.
Removing SGM-based illumination conditioning leads to unstable shading estimates and increased illumination leakage into albedo.
@@ 4.7 Qualitative Results
Figure 4 shows qualitative results on Hypersim and InteriorVerse. IDT produces sharper depth, cleaner normals, and more coherent albedo and shading across views. Compared to baselines, IDT better isolates non-Lambertian effects and maintains consistent intrinsic factors under viewpoint changes.
@@ 5 Conclusion
Limitations. IDT relies on a simplified image formation model and may be challenged by extreme lighting conditions or strongly non-Lambertian materials. In addition, the current approach does not explicitly enforce global geometric constraints and is primarily evaluated on indoor scenes.
We introduced Intrinsic Decomposition Transformer (IDT), a feed-forward framework for physically grounded multi-view intrinsic image decomposition. By jointly reasoning over multiple views with transformer attention, IDT infers view-consistent intrinsic factors in a single forward pass without iterative inference. An explicit image formation model that separates diffuse reflectance, diffuse shading, and specular effects further improves the disentanglement of material and illumination. Experiments on synthetic and real-world indoor datasets demonstrate that IDT produces cleaner intrinsic decompositions and substantially improved multi-view consistency over prior methods.